What is the [H+] in 0.025 M NaOH?
1
Expert's answer
2020-03-27T09:00:51-0400

Solution:

NaOH is a strong base. This means it completely dissociates into its ions.

NaOH = Na+ + OH-

Therefore, the concentration of OH is 0.025M.

[OH-] = 0.025.


1) Therefore, we can calculate the pOH:

pOH = −log([OH]) = −log(0.025) = 1.602

Therefore, now we can calculate the pH:

pH = 14 − pOH = 14 − 1.602 = 12.398

Now we can calculate the concentration of H+:

pH = -log([H+]);

[H+] = 10- pH = 10-12.398 = 3.99*10-13 = 4*10-13

[H+] = 4*10-13 M


or


2) pH + pOH = 14;

-log[H+] + (-log[OH-]) = 14;

[H+] * [OH-] = 10-14;

[H+] = 10-14 / [OH-];

[H+] = 10-14 / 0.025 = 4*10-13

[H+] = 4*10-13 M


Answer: 4*10-13 M is the [H+] in 0.025 M NaOH.